OpenStudy (anonymous):

Simplify the following expression. 5y(7y - 6) + 27y 35y2 - 3y 35y2 + 27y - 6 97y - 6 35y2 + 3y

OpenStudy (richyw):

\[5y(7y-6)+27y\]\[=35y^2-30y+27y\]\[=35y^2-3y\]\[=y(35y-3)\]

OpenStudy (anonymous):

Alright, you would be exactly correct except for your +3, in your work, when you exapdn 5y(7y-6) you get 35y^2-30y. Then all you have to do is add +25 to -30. Then you get 35y^2-3